# On A Scale From 1 to 5: Quantifying Hallucination in Faithfulness Evaluation

### v1

@@ 1 Introduction
We investigated two approaches, namely using LLMs with a grading rubrics and natural language inference (NLI) models, to derive a faithfulness score for a given reference and hypothesis pair. We experimented with 5 LLMs and 3 variants of NLI models on 4 travel-domain dataset. We studied the score variation by comparing various types of hypothesis: gold hypothesis, sentence-level gold hypothesis, intrinsic hypothesis, and extrinsic hypothesis. In addition, we explored the model sensitivity in picking up different levels of hallucination. Our main contributions are: First, we showed that LLMs can accurately distinguish faithfulness apart from hallucination, and that reasoning during grading help LLMs gain sensitivity towards hallucination. Second, we presented methods to generate synthetic hallucination by type. Third, we demonstrated tuning with synthetic data can improve NLI model’s performance. Lastly, we illustrated the progression of scores based on the percentage of hallucinating content.
@@ 3 System
We evaluated faithfulness scoring models on 4 travel-domain industry datasets. It should be noted that our dataset are taken from production and only contains gold data. We discuss means to create synthetic hallucination dataset in Section 3.1. Statistics of the dataset can be found in Table 1.
@@ 4 Result & Discussion
In this section, we discuss faithfulness scoring on the reference and hypothesis pairs for two sets of experiments. The score ranges from "1-very unfaithful" to "5-very faithful", with discrete intervals "2-very faithful", "3-slightly unfaithful", and "4-very unfaithful". It should be noted that HHEM models return a continuous probabilistic value in the range [0, 1]. It was later scaled to [1, 5] for ease of interpretation.
@@ 4.1 Scoring Quality
HHEM models are also less sensitive toward extrinsic hallucinations as compared to intrinsic hallucinations. Fine-tuning on extrinsic synthetic JSON data alone helped HHEM-extrinsic model gain better performance on almost all tasks. However, during fine-tuning, we noticed that it was rather easy to over-tune the HHEM model and made it less tolerate to any type of creative contents. We first tried tuning with 400 each gold and hallucination examples for 5 epochs, which resulted in overly strict scores. By reducing the number of epochs to 3, we were able to achieve the performance as shown in Figure 2. We believe if the model is trained on more diverse synthetic examples with a carefully crafted dataset, the results would be more stable than what we have shown here.
@@ 6 Limitations
    Due to resource limitations, we focused on testing only travel-domain dataset, but did not include popular open source faithfulness dataset. We justify that real-world dataset are often more challenging than research domain dataset, as there exist more noise in real-world data. We believe our approach can be generalizable to other domains.
@@
    Due to legal and privacy constraints, we do not plan to release our dataset as some of it contains sensitive customer information.

### v2

@@ 1 Introduction
We investigated two approaches, namely using LLMs with a grading rubrics and natural language inference (NLI) models, to derive a faithfulness score for a given reference and hypothesis pair. We experimented with 5 LLMs and 3 variants of NLI models on 4 travel-domain dataset. We studied the score variation by comparing various types of hypothesis: gold hypothesis, sentence-level gold hypothesis, intrinsic hypothesis, and extrinsic hypothesis. In addition, we explored the model sensitivity in picking up different levels of hallucination. Our main contributions are: First, we showed that LLMs can accurately distinguish faithfulness apart from hallucination, and that reasoning during grading help LLMs gain sensitivity towards hallucination. Second, we presented methods to generate synthetic hallucination by type and demonstrated that tuning with synthetic data can improve NLI model’s performance. Lastly, we illustrated the progression of scores based on the percentage of hallucinating content.
@@ 3 System
We evaluated faithfulness scoring models on 4 travel-domain industry datasets. It should be noted that our datasets are taken from production and only contains gold data. We discuss means to create synthetic hallucination dataset in Section 3.1. Statistics of the dataset can be found in Table 1.
@@ 4 Result & Discussion
In this section, we discuss faithfulness scoring on the reference and hypothesis pairs for two sets of experiments. The score ranges from "1-highly unfaithful", "2-very faithful", "3-slightly unfaithful", and "4-very unfaithful", to "5-highly faithful". It should be noted that HHEM models return a continuous probabilistic value in the range [0, 1]. It was later scaled to [1, 5] for ease of interpretation.
@@ 4.1 Scoring Quality
HHEM models are also less sensitive toward extrinsic hallucinations as compared to intrinsic hallucinations. Fine-tuning on extrinsic synthetic JSON data alone helped HHEM-extrinsic model gain better performance on almost all tasks. However, during fine-tuning, we noticed that it was rather easy to over-tune the HHEM model and made it less tolerant to any type of creative contents. We first tried tuning with 400 each gold and hallucination examples for 5 epochs, which resulted in overly strict scores. By reducing the number of epochs to 3, we were able to achieve the performance as shown in Figure 2. We believe if the model is trained on more diverse synthetic examples with a carefully crafted dataset, the results would be more stable than what we have shown here.
@@ 6 Limitations
    Due to resource limitations, we focused on testing only travel-domain dataset, but did not include popular open source faithfulness dataset. We justify that real-world dataset are often more challenging than research domain dataset, as there exist more noise in real-world data. We believe our approach can be generalized to other domains.
@@
    Due to privacy concerns, we plan to release a limited portion of the dataset. We’re in the process of legal review to determine which sections of the dataset can be made public.
@@ A.3 Additional Results
A.4 Additional Discussion on Evaluation Subjectivity
One may argue that a rubric-based approach is subject to to the particular use case which can be challenging across different contexts. This concern stands for not only the method proposed in this paper, but also other polarity-based faithfulness classification models. Similar to any traditional machine learning methods, the definition or gold data is defined by human annotators. A model can also be over-fitting to a particular use case if the gold labels are very domain specific. For example, in opinion mining, "A is very fast" can indicate positive opinion if "A" is a calculator, but it can also indicate negative opinion if "A" is a person driving in the resident area full of kids. Now the example might seem a bit extreme, the point we’re trying to make is that while consistency is very important, subjectivity is not 100% bad and it can sometimes help domain-specific adaptions.
In our case, the subjectivity only resides in the quantifiable metrics, but not in the domain itself (e.g. travel). In Table 5, we attached the prompt used to score summaries during the experiments. The prompt is domain independent which can be applied to any summarization tasks. Because some quantifiable metrics such as adjective usage and tone of voice can be subjective depending on the use case, having the ability for the user to tailor the rubrics will best help the LLM judge fit to that particular use case. For example, it is possible in stylized text generation that having LLMs add descriptive adjectives and verbs is desired in the generation, in which case, the user can adjust the “adjective consistency” requirements in the rubrics to fit that use case. We intended to introduce an adaptable solution, and similar to any other ML models, small adjustment might still be required when adapting to granular details.
A.5 Additional Discussion on Organic vs. Synthetic Hallucinations
Due to the lack of organic hallucinations in applications, we focused on synthetic hallucinations generated based on real-world observations. One may raise the concern that it is risky to draw conclusions on synthetic hallucinations over organic ones, as the distribution may not be fully modeled. We agree that organic hallucinations are more valuable and that in ideal scenarios, we should develop such faithfulness models using orgnic data as much as we can. However, we believe in the real-world case by case scenario, it is rather difficult to obtain a large portion of organic hallucinations for development. We believe for a well-defined system with a trained developer, even the initial development stage would suffer from insufficient negative examples. However, this does not mean the evaluation should be omitted. It is rather common for developer to generate synthetic test cases (or edge cases) for testing. Thus, we tailored the synthetic data generation for each use case to tackle tricky and subtle inconsistencies.
We believe the conclusion on the experiments can still be generalized because 1) Real-world dataset often contains more noise and are more complex in nature than academic gold dataset. For example, ConvoTS datasets contains many turns of redundant user interactions due to the limitation of the virtual assistant. 2) A portion of the synthetic changes are tricky enough to mimic organic hallucination (see examples in Table 7). Furthermore, even the best LLM judge (GPT-4) failed on some of those tricky cases.